OpenStudy (anonymous):

PLEASE HELP Me i need to know this to do my LAST assignments so i can get back into public school. How can the properties of logarithms be applied to simplify or solve equations?

OpenStudy (anonymous):

@Abhisar please can u help me!!!????

OpenStudy (anonymous):

depending on properties... if you have m and n for example as equations... you want to get their product such as m x n... we can use \[\log mn = \log m + \log n\]... it is simply addition of log of each equation...to summarize other properties \[\log\frac{m}{n}=\log m-\log n\]\[\log m^n=n\log m\]\[\log{\sqrt[n]{m}}=\frac{\log m}{n}\]